OpenStudy (anonymous):

Find the solution to the equations. 3x - y = -4 x + y = 0

OpenStudy (anonymous):

x=-2 y= -2

OpenStudy (help!!!!):

Just add up the equations

OpenStudy (anonymous):

we have to add both the equations 3x-y+x+y=-4+0

OpenStudy (anonymous):

here,y gets cancelled.then we get, 3x+x=-4 4x=-4 x=-4/4 x=-1

OpenStudy (anonymous):

........................

OpenStudy (anonymous):

we got the value of x, now put the value of x in any one of the above equations,say, x+y=0 here put the value of x and get y. hence we get both x and y.

OpenStudy (anonymous):

those are the answers it gives me to pick from A.(0, 0) B.(-1, 1) C.(1, -1)

OpenStudy (anonymous):

its B right?